# Early stopping for conjugate gradients in statistical inverse problems

### v1

@@ 5.1 Decomposition of the prediction error
We deduce by direct calculation
@@
When restricting to singular values $\lambda_{i}^{2}\leqslant x_{1,t}$ the same calculation gives
@@
On the other hand, we have $\lVert A(\widehat{f}_{t}-f)\rVert_{>}^{2}=\lVert(\mathsf{1-r_{t}})Y-g\rVert_{>}^{2}\leqslant 2(\lVert\xi\rVert_{>}^{2}+\lVert\mathsf{r_{t}}Y\rVert_{>}^{2})$ and thus
$\tfrac{1}{2}\lVert A(\widehat{f}_{t}-f)\rVert^{2}$ $\leqslant\lVert A(\widehat{f}_{t}-f)\rVert_{<}^{2}+\tfrac{1}{2}\lVert A(\widehat{f}_{t}-f)\rVert_{>}^{2}$
$\leqslant\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}+\lVert\mathsf{r_{t}}Y\rVert_{<}^{2}-\lVert\mathsf{r_{t,<}^{1/2}}Y\rVert^{2}+\lVert\mathsf{(1-r_{t})_{<}^{1/2}}\xi\rVert^{2}+\lVert\xi\rVert_{>}^{2}+\lVert\mathsf{r_{t}}Y\rVert_{>}^{2}$
@@
[…]
@@
    $S_{t,\lambda}\leqslant\lVert((\mathsf{\lvert r_{t}^{\prime}(0)\rvert x)^{1/2}\wedge 1})\xi\rVert^{2}$.
[…]
  1. (g)
@@
  2. (h)
    the upper bound $\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}$ is nonincreasing in $t$.
@@
For later purposes we also need the following relatively rough bound. The proof of Lemma 5.5 can be found in Section 9.4.
@@ 5.2 An oracle minimax bound for the prediction error
with a constant $C_{\mu,p,C_{A}}>0$ depending only on $\mu$, $p$ and $C_{A}$.
@@
By 5.3(g) and 4.8, the approximation error satisfies
@@
Combining the bound in 5.3(d) and (5.4) with Proposition 5.2 yields the first claim.
Assume (GN) and (PSD($p$,00,$C_{A}$)) with $p>1/2$. Since the upper bound for $A_{t,\lambda}$ is decreasing in $\lvert r_{t}^{\prime}(0)\rvert$ and the upper bound for $S_{t,\lambda}$ is increasing in $\lvert r_{t}^{\prime}(0)\rvert$ and $A_{\tau_{\mathfrak{b}},\lambda}=S_{\tau_{\mathfrak{b}},\lambda}$, we obtain by Proposition 5.2
@@
$\leqslant 4\inf_{\rho>0}\Big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu-1}+C_{A}^{2}\delta^{2}\sum_{i=1}^{D}\big{(}(\rho i^{-2p})\wedge 1\big{)}\Big{)}.$
@@
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\big{]}\leqslant 4\inf_{\rho>0}\big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu-1}+C_{A}^{2}\delta^{2}\tfrac{2p}{2p-1}\rho^{1/(2p)}\big{)}.$
@@
for $\delta/R\leqslant d_{\mu,p,c_{A}}$, provided $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$, where the infimum is taken over all ${\mathbb{R}}^{P}$-valued estimators $\widehat{f}$. The positive constants $c_{\mu,p,c_{A}}$, $d_{\mu,p,c_{A}}$ depend only on $\mu$, $p$ and $c_{A}$.
Remark 5.12.
@@
SVD entries of the signal such that we assume that $D$ is at least of this order. In this case, the error rate $\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$ is minimax.
In the case $D\ll t_{\mu,p,R}^{\mathrm{pred}}(\delta)$, we have $\operatorname{{\mathbb{E}}}[\lVert\xi\rVert^{2}]=\delta^{2}D\ll\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$ and thus by Remark 5.7 also $\operatorname{{\mathbb{E}}}[\lVert A(\widehat{f}_{\tau_{b}}-f)\rVert^{2}]\lesssim\delta^{2}D\ll\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$. In fact, the dimension is so small that regularisation is not needed and in a statistical sense the problem becomes parametric.
@@
The CG algorithm can be compared to Showalter’s method (or asymptotic regularisation), which is just the classical gradient flow and appears as continuous-time limit of standard iterative linear regularisation schemes. The Showalter estimator of $f$ at $s\geqslant 0$ is given by $\widehat{f}_{s}^{\mathrm{SW}}=A^{\dagger}(\mathsf{1-\exp(-sx)})Y$ [8, Example 4.7]. Its prediction error can be decomposed as
@@
Identifying $s\triangleq\tfrac{1}{2}\lvert r_{t}^{\prime}(0)\rvert$, the bias term $B_{s,\lambda}^{2}(f)$ coincides with the upper bound on $A_{t,\lambda}$ given in 5.3(g). Using that $1-e^{-y}\leqslant y\wedge 1\leqslant\tfrac{e}{e-1}(1-e^{-y})$ for $y\geqslant 0$, the stochastic error term $\lVert(\mathsf{1-e^{-sx}})\xi\rVert^{2}$ of $\widehat{f}_{s}^{\mathrm{SW}}$ is up to constants equal to $\lVert(\mathsf{sx\wedge 1})\xi\rVert^{2}$. For $s\triangleq\frac{1}{2}\lvert r_{t}^{\prime}(0)\rvert$, this is smaller than the upper bound $\lVert(\mathsf{(\lvert r_{t}^{\prime}(0)\rvert x)^{1/2}\wedge 1})\xi\rVert^{2}$ on $S_{t,\lambda}$ from 5.3(d).
We achieve, however, the same order of the oracle prediction error under mild conditions on the singular value decay. Assuming distinct singular values $(\lambda_{j})$ and
$\sum_{l=j}^{D}{\lambda}_{l}^{2}\lesssim j\lambda_{j}^{2},\quad j=2,\dots,D,$ (5.8)
we obtain for $\rho^{-1}\in({\lambda}_{j}^{2},{\lambda}_{j-1}^{2}]$
$\operatorname{{\mathbb{E}}}\big{[}\lVert(\mathsf{(\rho x)^{1/2}\wedge 1})\xi\rVert^{2}\big{]}=\delta^{2}\Big{(}\#\{i=1,\dots,D\,|\,\lambda_{i}^{2}\geqslant\rho^{-1}\}+\sum_{i:\,\lambda_{i}^{2}<\rho^{-1}}\rho\lambda_{i}^{2}\Big{)}$
$\leqslant\delta^{2}\Big{(}(j-1)+{\lambda}_{j}^{-2}\sum_{l=j}^{D}{\lambda}_{l}^{2}\Big{)}\lesssim\delta^{2}(j-1)=\operatorname{{\mathbb{E}}}\big{[}\lVert{\bf 1}(\mathsf{\rho x\geqslant 1})\xi\rVert^{2}\big{]}\leqslant\operatorname{{\mathbb{E}}}\big{[}\lVert(\mathsf{(\rho x)\wedge 1})\xi\rVert^{2}\big{]}.$ (5.9)
Combining all possible intervals for $\rho^{-1}$ and considering the case $\rho^{-1}\leqslant{\lambda}_{D}^{2}$ separately, the inequality (5.9) holds for all $\rho\geqslant\lambda_{1}^{-2}$. If $\tau_{\mathfrak{b}}\geqslant 1$, we have $\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert\geqslant x_{1,1}^{-1}\geqslant\lambda_{1}^{-2}$ by 4.5(g) and we can apply (5.9) for $\rho\triangleq\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert$. We then obtain by the monotonicity argument with respect to $\rho$ in the proof of Theorem 5.8 and the bias-variance decomposition (5.7) that
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\big{]}\leqslant 2\operatorname{{\mathbb{E}}}[A_{\tau_{\mathfrak{b}},\lambda}+S_{\tau_{\mathfrak{b}},\lambda}]\lesssim\inf_{s\geqslant\lambda_{1}^{-2}}\big{(}B_{s,\lambda}^{2}(f)+V_{s,\lambda}\big{)}=\inf_{s\geqslant\lambda_{1}^{-2}}\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{s}^{\mathrm{SW}}-f)\rVert^{2}\big{]}.$
It is not difficult to see that for sufficiently small $\delta>0$ we will have $\tau_{\mathfrak{b}}\geqslant 1$ and $\operatorname*{argmin}_{s\geqslant 0}\operatorname{{\mathbb{E}}}[\lVert A(\widehat{f}_{s}^{\mathrm{SW}}-f)\rVert^{2}]\geqslant\lambda_{1}^{-2}$ whenever the signal $f$ is not trivial in the sense that at least two SVD entries $f_{i}$ are nonzero.
@@
The singular value condition (5.8) is easily verified under polynomial singular value decay (PSD($p$,$c_{A}$,$C_{A}$)) with $p>1/2$, but is also true in the case of exponential or Gaussian-type decay. Specifically, for exponential decay $\lambda_{i}\thicksim e^{-pi}$, $p>0$, and polynomial source condition we obtain the prediction error rate
@@ 6 Early stopping and its prediction error
By the definition of $A_{t,\lambda}$, for any $t\in[0,d]$, we deduce
@@
This alternative expression of $\tau_{\mathfrak{b}}$ motivates the choice of the residual-based stopping rule $\tau$, as explained in Remark 6.2. Note that $R_{\tau}^{2}=\kappa$ and $R_{\tau_{\mathfrak{b}}}^{2}=\lVert\xi\rVert^{2}+2\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle$ hold by the continuity of $R_{t}^{2}$ in $t$.
@@
Lemma 4.3 implies
@@
$=\operatorname{{\mathbb{E}}}\big{[}\lvert\lVert\xi\rVert^{2}+2\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle-\kappa\rvert\big{]}$
@@
Due to $\operatorname{{\mathbb{E}}}[\lVert\xi\rVert^{2}]^{1/2}=\delta\sqrt{D}$, this is unfortunately highly suboptimal under (GN). If $r_{\tau_{\mathfrak{b}},<}$ were independent of $\xi$ or even deterministic as for the cross term in linear methods [3, 4], then the inner product would have expectation zero and standard deviation $\delta\operatorname{{\mathbb{E}}}[\lVert\mathsf{r_{\tau_{\mathfrak{b}},<}}g\rVert^{2}]^{1/2}$, a dimensionless and much tighter bound than (6.2). In our case of intricate randomness of $r_{\tau_{\mathfrak{b}},<}$, we are going to show a slightly larger bound that grows like $\sqrt{\log d}$ in the dimension, but still establishes a sufficiently low order of the cross term, see Proposition 6.8. The probabilistic proof uses concentration of self-normalised Gaussian processes by analysing the potential complexity of $\mathsf{r_{\tau_{\mathfrak{b}},<}}g$ in ${\mathbb{R}}^{D}$. Before, a high probability upper bound on $\lvert r_{t}^{\prime}(0)\rvert$ will be established, giving in Corollary 6.6 an at least linear growth of $t\mapsto S_{t,\lambda}$.
In the sequel, we sometimes write $\lambda(\boldsymbol{\cdot})^{-2}$ instead of $\lambda_{\boldsymbol{\cdot}}^{-2}$ for notational convenience.
@@
Under (GN), we have for all $z>0$ with probability at least $1-e^{-z}$
$\forall\,t\in[0,d]:\;\lvert r_{t}^{\prime}(0)\rvert\wedge\lambda_{D}^{-2}\leqslant\lambda\Big{(}\big{(}\lfloor\tfrac{2e}{e-1}\delta^{-2}S_{t,\lambda}+4(z+\log d)\rfloor+1\big{)}\wedge D\Big{)}^{-2}.$
@@
$S_{t,\lambda}=\lVert\mathsf{(1-r_{t,<})^{1/2}}\xi\rVert^{2}\geqslant\lVert\mathsf{(1-\exp(-\lvert r_{t}^{\prime}(0)\rvert x))^{1/2}}\xi\rVert^{2}\geqslant(1-e^{-1})\sum_{i:\,\lambda_{i}^{-2}\leqslant\lvert r_{t}^{\prime}(0)\rvert}\xi_{i}^{2}.$
@@
On $\bigcap_{\rho>0}\Omega(\rho)\subset\Omega(\lvert r_{t}^{\prime}(0)\rvert)$, this gives

### v2

@@ 5.1 Decomposition of the prediction error
For $t=0$, the claim is clear, since $r_{0,<}\equiv 1$, $r_{0,>}\equiv 0$ and $\widehat{f}_{0}=0$. Let $t>0$. We deduce by direct calculation
@@
When restricting to singular values $\lambda_{i}^{2}<x_{1,t}$ the same calculation gives
@@
On the other hand, we have $\lVert A(\widehat{f}_{t}-f)\rVert_{\geqslant}^{2}=\lVert(\mathsf{1-r_{t}})Y-g\rVert_{\geqslant}^{2}\leqslant 2(\lVert\xi\rVert_{\geqslant}^{2}+\lVert\mathsf{r_{t}}Y\rVert_{>}^{2})$ and thus
$\tfrac{1}{2}\lVert A(\widehat{f}_{t}-f)\rVert^{2}$ $\leqslant\lVert A(\widehat{f}_{t}-f)\rVert_{<}^{2}+\tfrac{1}{2}\lVert A(\widehat{f}_{t}-f)\rVert_{\geqslant}^{2}$
$\leqslant\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}+\lVert\mathsf{r_{t}}Y\rVert_{<}^{2}-\lVert\mathsf{r_{t,<}^{1/2}}Y\rVert^{2}+\lVert\mathsf{(1-r_{t})_{<}^{1/2}}\xi\rVert^{2}+\lVert\xi\rVert_{\geqslant}^{2}+\lVert\mathsf{r_{t}}Y\rVert_{>}^{2}$
@@
The stochastic error term $S_{t,\lambda}$, $t\in[0,d]$, satisfies
  1. (a)
    $S_{t,\lambda}=0$ for $\xi=0$,
  2. (b)
    $S_{0,\lambda}=0$ and $S_{d,\lambda}=\lVert\xi\rVert^{2}$,
  3. (c)
    $S_{t,\lambda}\leqslant\lVert((\mathsf{\lvert r_{t}^{\prime}(0)\rvert x)^{1/2}\wedge 1})\xi\rVert^{2}$,
@@
    $S_{t,\lambda}$ is nondecreasing in $t$,
  5. (e)
    $S_{t,\lambda}$ is continuous in $t$.
The approximation error term $A_{t,\lambda}$, $t\in[0,d]$, satisfies
  1. ()
    $A_{t,\lambda}\leqslant 0$ if $Af=0$,
  2. (a)
    $A_{0,\lambda}=\lVert Af\rVert^{2}$ and $A_{d,\lambda}=0$,
  3. (b)
@@
  4. (c)
    the upper bound $\lVert\mathsf{r_{t,<}^{1/2}}g\rVert^{2}$ is nonincreasing in $t$,
  5. (d)
    $A_{t,\lambda}$ is continuous in $t$.
@@
For later purposes we also need the following relatively rough bound, whose proof can be found in Section 9.4.
@@ 5.2 An oracle minimax bound for the prediction error
with $\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$ given in (3.2) and a constant $C_{\mu,p,C_{A}}>0$ depending only on $\mu$, $p$ and $C_{A}$.
@@
By 5.3(b) and 4.8, the approximation error satisfies
@@
Combining the bound in 5.3(c) and (5.4) with Proposition 5.2 yields the first claim.
Assume (GN) and (PSD($p$,00,$C_{A}$)) with $p>1/2$. If $Af=0$, the second claim is clear. Suppose $Af\neq 0$ such that $\tau_{\mathfrak{b}}>0$ and $\lvert r_{\tau_{\mathfrak{b}}}^{\prime}(0)\rvert>0$. Since the upper bound for $A_{t,\lambda}$ is decreasing in $\lvert r_{t}^{\prime}(0)\rvert$ and the upper bound for $S_{t,\lambda}$ is increasing in $\lvert r_{t}^{\prime}(0)\rvert$ and $A_{\tau_{\mathfrak{b}},\lambda}=S_{\tau_{\mathfrak{b}},\lambda}$, we obtain by Proposition 5.2
@@
$\leqslant 4\inf_{\rho>0}\Big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu-1}+(C_{A}\vee 1)^{2}\delta^{2}\sum_{i=1}^{D}\big{(}(\rho i^{-2p})\wedge 1\big{)}\Big{)}.$
@@
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\big{]}\leqslant 4\inf_{\rho>0}\big{(}R^{2}(\mu+1/2)^{2\mu+1}\rho^{-2\mu-1}+(C_{A}\vee 1)^{2}\delta^{2}\tfrac{2p}{2p-1}\rho^{1/(2p)}\big{)}.$
@@
for $\delta/R\leqslant\tilde{c}_{\mu,p,c_{A}}$, provided $(R^{2}\delta^{-2})^{1/(4\mu p+2p+1)}\lesssim_{\mu,p,c_{A}}D$, where the infimum is taken over all ${\mathbb{R}}^{P}$-valued estimators $\widehat{f}$ and $\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$ is given in (3.2). The positive constants $c_{\mu,p,c_{A}}$, $\tilde{c}_{\mu,p,c_{A}}$ depend only on $\mu$, $p$ and $c_{A}$.
Remark 5.12 (Role of the dimension).
@@
SVD entries of the signal such that we assume that $D$ is at least of this order. In this case, the error rate $\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$ is minimax. In the case $D\ll t_{\mu,p,R}^{\mathrm{pred}}(\delta)$, we have $\operatorname{{\mathbb{E}}}[\lVert A(\widehat{f}_{d}-f)\rVert^{2}]=\operatorname{{\mathbb{E}}}[\lVert\xi\rVert^{2}]=\delta^{2}D\ll\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$ and thus by Remark 5.7 also $\operatorname{{\mathbb{E}}}[\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}]\lesssim\delta^{2}D\ll\mathcal{R}^{*,\mathrm{pred}}_{\mu,p,R}(\delta)$. In fact, the dimension is so small that regularisation is not needed and iterating until the final index $d$ is optimal. The corresponding lower bound for the prediction error is indeed of order $\delta^{2}D$, which can be shown by the generalised Fano method.
If we view (2.1) as a discretisation of an underlying continuous inverse problem, which we approach as $D(\delta),P(\delta)\to\infty$ for $\delta\to 0$, we might rather want to consider sequences $f^{(\delta)}\in{\mathbb{R}}^{P(\delta)}$, $A^{(\delta)}\in{\mathbb{R}}^{D(\delta)\times P(\delta)}$ of discretised functions and operators satisfying the source condition (2.2) and the spectral decay condition (PSD($p_{\delta}$,$c_{A,\delta}$,$C_{A,\delta}$)) with possibly varying parameters $\mu_{\delta},R_{\delta},p_{\delta},c_{A,\delta},C_{A,\delta}$. If these parameters remain bounded away from zero and infinity and $p_{\delta}>1/2+\varepsilon$ for some fixed $\varepsilon>0$, then the same proofs, but with much heavier notation, yield in Theorem 5.8
$\limsup_{\delta\to 0}\mathcal{R}^{*,\mathrm{pred}}_{\mu_{\delta},p_{\delta},R_{\delta}}(\delta)^{-1}\operatorname{{\mathbb{E}}}\big{[}\lVert A^{(\delta)}(\widehat{f}_{\tau_{\mathfrak{b}}}-f^{(\delta)})\rVert^{2}\big{]}<\infty$
and in Proposition 5.11
$\liminf_{\delta\to 0}\mathcal{R}^{*,\mathrm{pred}}_{\mu_{\delta},p_{\delta},R_{\delta}}(\delta)^{-1}\inf_{\widehat{f}}\sup_{f:\,\lVert f\rVert_{\mu_{\delta}}\leqslant R_{\delta}}\operatorname{{\mathbb{E}}}_{f}\big{[}\lVert A^{(\delta)}(\widehat{f}-f)\rVert^{2}\big{]}>0,$
provided $\lim_{\delta\to 0}D(\delta)\delta^{2/(4\mu_{\delta}p_{\delta}+2p_{\delta}+1)}=\infty$. Similar extensions also hold for all subsequent results.
@@
The CG algorithm can be compared to Showalter’s method (or asymptotic regularisation), which is just the classical gradient flow and appears as continuous-time limit of standard iterative linear regularisation schemes. The Showalter estimator of $f$ at $s\geqslant 0$ is given by $\widehat{f}_{s}^{\mathrm{SW}}=A^{\dagger}(\mathsf{1-\exp(-sx)})Y$ [8, Example 4.7]. Its prediction error for deterministic $s\geqslant 0$ can be decomposed as
@@
Inserting $s\triangleq\tfrac{1}{2}\lvert r_{t}^{\prime}(0)\rvert$, the then random term $\lVert\mathsf{e^{-sx}}Af\rVert^{2}$ coincides with the upper bound on $A_{t,\lambda}$ given in 5.3(b). Using that $1-e^{-y}\leqslant y\wedge 1\leqslant\tfrac{e}{e-1}(1-e^{-y})$ for $y\geqslant 0$, the stochastic error term $\lVert(\mathsf{1-e^{-sx}})\xi\rVert^{2}$ of $\widehat{f}_{s}^{\mathrm{SW}}$ is up to constants equal to $\lVert(\mathsf{sx\wedge 1})\xi\rVert^{2}$. For $s\triangleq\frac{1}{2}\lvert r_{t}^{\prime}(0)\rvert$, this is smaller than the upper bound $\lVert(\mathsf{(\lvert r_{t}^{\prime}(0)\rvert x)^{1/2}\wedge 1})\xi\rVert^{2}$ on $S_{t,\lambda}$ from 5.3(c).
We achieve, however, the same order of the oracle prediction error under the following mild condition on the singular value decay, where $C_{\lambda}>0$ is a constant:
$\sum_{i=j}^{D}{\lambda}_{i}^{2}\leqslant C_{\lambda}j\lambda_{j}^{2},\quad j=2,\dots,D.$ (5.8)
The condition (5.8) is easily verified under polynomial singular value decay (PSD($p$,$c_{A}$,$C_{A}$)) with $p>1/2$, but is also true in the case of exponential or Gaussian-type decay.
Assuming for simplicity distinct singular values $(\lambda_{j})$, we obtain for deterministic $\rho^{-1}\in({\lambda}_{j}^{2},{\lambda}_{j-1}^{2}]$
$\operatorname{{\mathbb{E}}}\big{[}\lVert(\mathsf{(\rho x)^{1/2}\wedge 1})\xi\rVert^{2}\big{]}=\delta^{2}\big{(}\#\{i=1,\dots,D\,|\,\lambda_{i}^{2}\geqslant\rho^{-1}\}+{\textstyle\sum_{i:\,\lambda_{i}^{2}<\rho^{-1}}}\rho\lambda_{i}^{2}\big{)}$
$\leqslant\delta^{2}\big{(}(j-1)+{\lambda}_{j}^{-2}\textstyle{\sum_{i=j}^{D}}{\lambda}_{i}^{2}\big{)}\leqslant(2C_{\lambda}+1)\delta^{2}(j-1)$
$=(2C_{\lambda}+1)\operatorname{{\mathbb{E}}}\big{[}\lVert{\bf 1}(\mathsf{\rho x\geqslant 1})\xi\rVert^{2}\big{]}\leqslant(2C_{\lambda}+1)\operatorname{{\mathbb{E}}}\big{[}\lVert(\mathsf{(\rho x)\wedge 1})\xi\rVert^{2}\big{]}.$ (5.9)
Combining all possible intervals for $\rho^{-1}$ and considering the case $\rho^{-1}\leqslant{\lambda}_{D}^{2}$ separately, the inequality (5.9) holds for all $\rho\geqslant\lambda_{1}^{-2}$. We then obtain by the monotonicity argument with respect to $\rho$ in the proof of Theorem 5.8 and the bias-variance decomposition (5.7) that
$\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{\tau_{\mathfrak{b}}}-f)\rVert^{2}\big{]}$ $\leqslant 2\operatorname{{\mathbb{E}}}[A_{\tau_{\mathfrak{b}},\lambda}+S_{\tau_{\mathfrak{b}},\lambda}]$
$\lesssim\inf_{\rho\geqslant\lambda_{1}^{-2}}\big{(}\lVert\mathsf{e^{-\rho x/2}}Af\rVert^{2}+\operatorname{{\mathbb{E}}}\big{[}\lVert(\mathsf{(\rho x)^{1/2}\wedge 1})\xi\rVert^{2}\big{]}\big{)}$
$\lesssim_{C_{\lambda}}\inf_{s\geqslant\lambda_{1}^{-2}}\big{(}B_{s,\lambda}^{2}(f)+V_{s,\lambda}\big{)}=\inf_{s\geqslant\lambda_{1}^{-2}}\operatorname{{\mathbb{E}}}\big{[}\lVert A(\widehat{f}_{s}^{\mathrm{SW}}-f)\rVert^{2}\big{]}.$
The infimum on the right-hand side can be taken over $s\geqslant 0$ for sufficiently small $\delta>0$ whenever $f\neq 0$.
@@
For exponential decay $\lambda_{i}\thicksim e^{-pi}$, $p>0$, and polynomial source condition, we obtain similarly to the proof of (5.3) the prediction error rate
@@ 6 Early stopping and its prediction error
By the definitions of $A_{t,\lambda}$ and $S_{t,\lambda}$, for any $t\in[0,d]$, we deduce
@@
This alternative expression of $\tau_{\mathfrak{b}}$ motivates the choice of the residual-based stopping rule $\tau$, as explained in Remark 6.2. Note that $R_{\tau_{\mathfrak{b}}}^{2}=\lVert\xi\rVert^{2}+2\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle$ holds due to $A_{\tau_{\mathfrak{b}},\lambda}=S_{\tau_{\mathfrak{b}},\lambda}$.
@@
In the case $\kappa\leqslant\lVert Y\rVert^{2}$, we have $R_{\tau}^{2}=\kappa$. Otherwise, if $\kappa>\lVert Y\rVert^{2}$, we stop immediately at $\tau=0$ such that $0\leqslant R_{\tau}^{2}-R_{\tau_{\mathfrak{b}}}^{2}=\lVert Y\rVert^{2}-R_{\tau_{\mathfrak{b}}}^{2}<\kappa-R_{\tau_{\mathfrak{b}}}^{2}$. Hence, we have in both cases the inequality
$\lvert R_{\tau}^{2}-R_{\tau_{\mathfrak{b}}}^{2}\rvert\leqslant\lvert\kappa-R_{\tau_{\mathfrak{b}}}^{2}\rvert.$
4.3(c) implies
@@
$\leqslant\operatorname{{\mathbb{E}}}\big{[}\lvert\kappa-\lVert\xi\rVert^{2}-2\langle{\xi},{\mathsf{r_{\tau_{\mathfrak{b}},<}}g}\rangle\rvert\big{]}$
@@
Due to $\operatorname{{\mathbb{E}}}[\lVert\xi\rVert^{2}]^{1/2}=\delta\sqrt{D}$, this is unfortunately highly suboptimal under (GN). If $r_{\tau_{\mathfrak{b}},<}$ were independent of $\xi$ or even deterministic as for the cross term in linear methods [3, 4], then the inner product would have expectation zero and standard deviation $\delta\operatorname{{\mathbb{E}}}[\lVert\mathsf{r_{\tau_{\mathfrak{b}},<}}g\rVert^{2}]^{1/2}$, a dimensionless and much tighter bound than (6.2).
In our case of intricate randomness of $r_{\tau_{\mathfrak{b}},<}$, we are going to show a slightly larger bound that grows like $\sqrt{\log d}$ in the dimension, but still establishes a sufficiently low order of the cross term, see Proposition 6.7. The probabilistic proof uses concentration of self-normalised Gaussian processes by analysing the potential complexity of $\mathsf{r_{\tau_{\mathfrak{b}},<}}g$ in ${\mathbb{R}}^{D}$. Before, a high probability upper bound on $\lvert r_{t}^{\prime}(0)\rvert$ will be established, giving in Corollary 6.5 an at least linear growth of $t\mapsto S_{t,\lambda}$.
In the sequel, we sometimes write $\lambda(i)^{-2}$ instead of $\lambda_{i}^{-2}$ for notational convenience, where $i\in\{1,\dots,D\}$.
@@
Under (GN), we have for all $z>0$ on an event of probability at least $1-e^{-z}$ the following two implications for all $t\in[0,d]$:
$\lvert r_{t}^{\prime}(0)\rvert<\lambda_{D}^{-2}\;\Rightarrow\;\lvert r_{t}^{\prime}(0)\rvert<\lambda\Big{(}\big{(}\lfloor\tfrac{2e}{e-1}\delta^{-2}S_{t,\lambda}+4(z+\log d)\rfloor+1\big{)}\wedge D\Big{)}^{-2},$
$\lvert r_{t}^{\prime}(0)\rvert\geqslant\lambda_{D}^{-2}\;\Rightarrow\;D\leqslant\lfloor\tfrac{2e}{e-1}\delta^{-2}S_{t,\lambda}+4(z+\log d)\rfloor.$
@@
$S_{t,\lambda}=\lVert\mathsf{(1-r_{t,<})^{1/2}}\xi\rVert^{2}\geqslant\lVert\mathsf{(1-\exp(-\lvert r_{t}^{\prime}(0)\rvert x))^{1/2}}\xi\rVert^{2}\geqslant(1-e^{-1})\sum_{i:\,\lambda_{i}^{-2}\leqslant\lvert r_{t}^{\prime}(0)\rvert}\xi_{i}^{2}.$ (6.3)
@@
On $\bigcap_{\rho>0}\Omega(\rho)\subset\Omega(\lvert r_{t}^{\prime}(0)\rvert)$, the definition of $\Omega(\lvert r_{t}^{\prime}(0)\rvert)$ and (6.3) give